One canned juice drink is 15% orange juice; another is 5% orange juice. How many liters of each should be mixed together in order to get 10L that is 10% orange juice?
Found 2 solutions by Alan3354, jorel1380:
Answer by Alan3354(69443)

10% is the average of 5% and 15%, so it's equal parts.
5 liters of each.

Answer by jorel1380(3719)

If you need 10% of 10L of orange juice, then you need 10X.1, or one liter of pure whole orange juice.
X+Y=10
.15X+.05Y=1
Multiply the second equation by 20, we get:
3X+Y=20
X+Y=10
2X=10
X=5
Y=5
We need 5L of each orange juice.